Divide and if possible simplify

ペ126 13 31.
Answer #1

261/3/131/3

= (26/13)1/3

= 21/3

If N= a*a*b then the cube root of N is roughly equal to the average of the three factors.

To find a cube root of 2, let us start with an estimate a=1.

2=1*1*2

Cube root of 2 is about (1+1+2)/3=4/3

Let us revise a to be 1.3.

2=1.3*1.3*1.18

Average of the three factors is 1.26.

A Cube root of 2 is 1.26 (approx)